Congratulations yet again. You've just been appointed chairman of the Council of Economic Advisers in Textland. You must rely on your research assistant for the specific numbers. He says income is $51,000, mpe is 0.6, and the president wants to raise output by $2,040. a. Advise him. (Increase/Decrease) government spending by $_________ or taxes by $_________. b. Your research assistant comes in and says "Sorry, I meant that the mpe is 0.4." You redo your calculations. (Increase/Decrease) government spending by $_________ or taxes by $_________. c. You're just about to see the president when your research assistant comes running in, saying, "Sorry, sorry, I meant that the mpe is 0.5." Redo your calculations. (Increase/Decrease) government spending by $_________ or taxes by $_________. a. Increase government spending by $816 or decrease taxes by $1360. b. Increase government spending by $1224 or decrease taxes by $3060. c. Increase government spending by $1020 or decrease taxes by $2040.
Answer #1

Ans.

a)  mpc = 0.6, required increase in output (Y) = $2040 , required increase in Government spending = G , required decrease in taxes = T

government spending multiplier (G) = 1/(1-mpc) = 1/(1-0.6) = 2.5

tax multiplier (T) = - mpc/(1-mpc) = -0.6/(1-0.6) = 1.5

Y = G*G 2040 = 2.5*G G = 2040/2.5 = 816

Y = T*T 2040 = 2.5*T T = 2040/2.5 = 1360

b)  mpc = 0.4, required increase in output (Y) = $2040 , required increase in Government spending = G , required decrease in taxes = T

government spending multiplier (G) = 1/(1-mpc) = 1/(1-0.4) = 1.67

tax multiplier (T) = - mpc/(1-mpc) = -0.4/(1-0.4) = -0.67

Y = G*G 2040 = 1.67*G G = 2040/1.67 = $1221.5

Y = T*T 2040 = 0.67*T T = 2040/0.67 = $3044.7

c)  mpc = 0.5, required increase in output (Y) = $2040 , required increase in Government spending = G , required decrease in taxes = T

government spending multiplier (G) = 1/(1-mpc) = 1/(1-0.5) = 2

tax multiplier (T) = - mpc/(1-mpc) = -0.5/(1-0.5) = -1

Y = G*G 2040 = 2*G G = 2040/2 = $1020

Y = T*T 2040 = 1*T T = 2040/1 = $2040